OpenStudy (anonymous):

Area of square is numerically less than twice its diagonal by 2. Find its Perimeter.

OpenStudy (anonymous):

area: x^2 diagonal: sqrt(2x^2) = x sqrt(2) 2x sqrt(2) - x^2 = 2

OpenStudy (radar):

Perimeter is 4x. Solve for x using quadratic formula. x=sqrt2 Perimeter =\[4\sqrt{2}\]